OpenStudy (anonymous):

PLEASE HELP A cylinder has a radius of 2 cm and a height of 5 cm. The surface area of the cylinder is _____ cm². 20 28 40 48

OpenStudy (anonymous):

The lateral area of a cylinder is the area of a rectangle whose dimensions are the height of the cylinder and the circumference of the base. The area of the bases is the sum of the areas of two circles with the radius of the base of the cylinder. A = 2(pi)rh + 2(pi)r^2 A = 2(3.14)(2)(5) + 2(3.14)(2^2) = ? You find the ? This looks like a test question I just had btw.

OpenStudy (anonymous):

87.92

OpenStudy (anonymous):

good :D

OpenStudy (anonymous):

thats not an option:(

OpenStudy (anonymous):

what are the options

OpenStudy (anonymous):

20 pie 28 pie 40 pie 48 pie

OpenStudy (anonymous):

i think it's 28

OpenStudy (anonymous):

ok to get it divide 87.92 by pi

OpenStudy (anonymous):

i was right!!!!!!!!!!!

OpenStudy (anonymous):

its 28

OpenStudy (anonymous):

20 x 3.14

OpenStudy (anonymous):

28****

OpenStudy (anonymous):

yeah its \[28*\pi \]